OpenStudy (kmmp):

Help please.. x^2=150

OpenStudy (kmmp):

I think I have it + -5 root 6

OpenStudy (anonymous):

close there is no +- jus simply \[5\sqrt{6}\]

OpenStudy (anonymous):

but great job :)

RhondaSommer (rhondasommer):

:)

zepdrix (zepdrix):

The plus/minus is correct :) You're taking the square root of a square, \(\rm \sqrt{x^2}=\pm x\)